Using the simultaneous equation method, find the values of x and y which satisfy the following equations:
3x+2y=22
2x+3y=23

Answer by alicealc(293)

multiply the first equation by 2:
6x + 4y = 44
multiply the second equation by 3:
6x + 9y = 69

use the elimination process,subtract the first equation by the second equation:
6x + 4y = 44
6x + 9y = 69
------------ (-)
-5y =-25
y = (-25)/(-5) = 5

substitute the y value to the first or second equation (here, I use the first):
6x + 4*5 =44
6x + 20= 44
6x = 44 - 20
6x = 24
x = 24/6 = 4